OpenStudy (anonymous):

How do you simplify things dealing with the laws of exponents, such as -3g^-2hk^-2/-6h^0

OpenStudy (anonymous):

idk sorry

OpenStudy (anonymous):

@pooja195 Would u know how to do this?

zepdrix (zepdrix):

\[\large\rm \frac{-3g^{-2}h k^{-2}}{-6h^0}\]Simplify this?

zepdrix (zepdrix):

This will be the most important rule that we can apply. \[\large\rm x^{-a}\quad=\frac{1}{x^a},\qquad\qquad\qquad \frac{1}{x^{-b}}\quad=\frac{x^b}{1}\] These give us ways to deal with negative exponents. The first one is the one we can make use of.

zepdrix (zepdrix):

\[\large\rm \frac{-3\color{orangered}{g^{-2}}h k^{-2}}{-6h^0}\]Let's see what happens to the g's,\[\large\rm \frac{-3h k^{-2}}{-6h^0\color{orangered}{g^{2}}}\]The exponent changes from negative to positive, and changes from numerator to denominator. We could put a 1 in it's place in the numerator, but notice that that's not necessary since other stuff is up there.

zepdrix (zepdrix):

Another important trick: \(\large\rm x^0=1\) If our base is being raised to the 0 power, we end up with a 1.

OpenStudy (anonymous):

Awesome! Thanks @zepdrix !